OpenStudy (anonymous):

Abcd is a square with a diagonal of 6 what is the length of a side and area of the square?

OpenStudy (anonymous):

S =d^2/2 = 6^2 : 2 = 18 a =b=c=d=\[\sqrt{18}\]

OpenStudy (anonymous):

S = d^2 :2 d=6 s= 6^2 : 2 = 18

OpenStudy (anonymous):

The side is \[\sqrt{18}\]

OpenStudy (anonymous):

Note the 45-45-90 triangle. That means the ratio of the legs to the hypotenuse (diagonal) is 1:sqrt(2) In this case, the legs will be 3sqrt(2), which means the area is 18.

jhonyy9 (jhonyy9):

- so than resulte that (diagonal)^2 = a^2 +b^2 but we know that a=b so than rtesult that (diagonal)^2 =2a^2 -so than 36=2a^2 - so a^2 = 18 - so a=sqrt18=3sqrt2 than a=b=3sqrt2 from this result that area of square willé be a multiply b what is =a squared so what is 3sqrt2 * 3sqrt2 = 18

jhonyy9 (jhonyy9):

ok ?